OpenStudy (anonymous):

1+cot^2(-theta) = csctheta I'm not allowed to change the right hand side of the equation How would I go about doing this problem?

OpenStudy (anonymous):

\[1+\cot ^2(-\theta)=\csc^2\theta \]

OpenStudy (michele_laino):

we can use this identity: \[\Large \cot \left( { - \theta } \right) = \frac{{\cos \left( { - \theta } \right)}}{{\sin \left( { - \theta } \right)}} = \frac{{\cos \theta }}{{ - \sin \theta }} = - \frac{{\cos \theta }}{{\sin \theta }}\]

OpenStudy (michele_laino):

@minisweet4

OpenStudy (anonymous):

How would that get me to \[\csc^2\theta\] in the end?

OpenStudy (isaiah.feynman):

You have to do some arithmetic to get there.

OpenStudy (michele_laino):

hint: if we apply that identity, we get: \[\large1 + {\left[ {\cot \left( { - \theta } \right)} \right]^2} = 1 + {\left( { - \frac{{\cos \theta }}{{\sin \theta }}} \right)^2} = 1 + \frac{{{{\left( {\cos \theta } \right)}^2}}}{{{{\left( {\sin \theta } \right)}^2}}}\]

OpenStudy (michele_laino):

please simplify this expression: \[\large 1 + {\left[ {\cot \left( { - \theta } \right)} \right]^2} = 1 + {\left( { - \frac{{\cos \theta }}{{\sin \theta }}} \right)^2} = 1 + \frac{{{{\left( {\cos \theta } \right)}^2}}}{{{{\left( {\sin \theta } \right)}^2}}} = ...?\]

OpenStudy (michele_laino):

furthermore, we can rewrite the right side of your expression, as below: \[\Large {\left( {\csc \theta } \right)^2} = \frac{1}{{{{\left( {\sin \theta } \right)}^2}}}\]

OpenStudy (anonymous):

I see what you did now

OpenStudy (isaiah.feynman):

That's very good!

OpenStudy (anonymous):

how'd you get rid of the negative in front of cos/sin?

OpenStudy (isaiah.feynman):

The square got rid of it.

OpenStudy (michele_laino):

since, we can write this: \[\large {\left( { - \frac{{\cos \theta }}{{\sin \theta }}} \right)^2} = \left( { - \frac{{\cos \theta }}{{\sin \theta }}} \right) \times \left( { - \frac{{\cos \theta }}{{\sin \theta }}} \right) = \frac{{{{\left( {\cos \theta } \right)}^2}}}{{{{\left( {\sin \theta } \right)}^2}}}\]

OpenStudy (anonymous):

Oh okay, thank you both very much. I may be putting up another problem after this one

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

Thank you!